Grade 12th passMechanics

A particle is projected vertically upwards from point A on the ground .It takes time t1 to reach point B but it still continues to move up.if it further takes t2 time to reach the ground from point B then height of point B from the ground is

Abhijeet
8 Years agoGrade 12th pass

1 Answer

Rituraj Tiwari
5 Years ago

The scenario you’re describing involves a particle that is projected vertically upwards, and we can determine the height of point B using the principles of kinematics. Let's break this down step by step, using the equations of motion that govern vertical motion under gravity.

Understanding Vertical Motion

When a particle is projected vertically upwards, it experiences a constant acceleration due to gravity, which we denote as g (approximately 9.81 m/s² downward). The motion can be analyzed in two parts: the upward journey to point B and the downward journey from point B to the ground.

Phase 1: Upward Motion to Point B

During the first phase, the particle reaches point B after time t1. The height h at point B can be calculated using the following kinematic equation:

  • h = u * t1 - 0.5 * g * t1²

Here, u is the initial velocity of the particle at point A. This equation accounts for the initial upward velocity and the effect of gravity pulling the particle back down.

Phase 2: Downward Motion from Point B

After reaching point B, the particle continues to move upwards until it reaches its maximum height, and then it starts to fall back down. The time taken to return to the ground from point B is t2. The total time for the upward journey to maximum height (from point B) and the downward journey back to the ground can be expressed as:

  • Total time to reach maximum height = t1 + t2
  • Time to fall from point B to the ground = t2

At the maximum height, the velocity becomes 0. We can use another kinematic equation to find the height attained from point B:

  • h = 0.5 * g * t2²

Final Height Calculation

To find the height of point B from the ground, we need to combine these equations. The height at point B can be expressed in terms of the total time taken to reach the maximum height:

  • h = u * t1 - 0.5 * g * t1²

For the downward motion from maximum height back to point B, we can rearrange our earlier findings to express the height from point B to the ground:

  • h = 0.5 * g * t2²

By equating these two expressions for height, we can solve for the height of point B:

  • u * t1 - 0.5 * g * t1² = 0.5 * g * t2²

Final Expression

From this equation, if we isolate the height h at point B, we can also express it in terms of time variables:

  • h = 0.5 * g * t2²

Thus, the height of point B can be effectively expressed in terms of the time it takes for the particle to fall back to the ground from that height, demonstrating the interplay between time and distance in vertical motion.

In summary, the height of point B from the ground can be determined using the time t2 it takes for the particle to fall back down, along with gravitational acceleration. This approach highlights the importance of understanding motion under constant acceleration, which is a fundamental principle in physics.